# Fusing Individualized Treatment Rules Using Secondary Outcomes: Supplementary Materials

### 19 Feb 2024 v2

@@
Appendix A ADDITIONAL DETAILS AND RESULTS FOR SECTIONS LABEL:sec:simulation AND LABEL:sec:realdata
In this section, we provide some details about implementation and experiment results in the simulation study of Section LABEL:sec:simulation and in the real data analysis of Section LABEL:sec:realdata.
@@ A.1 Implementation Details
A.2 Additional Details for Data Generation and Value Function Estimation in Section LABEL:sec:simulation
@@
A.3 Additional Results for Estimating FITR in Section LABEL:sec:simulation
@@ A.5 Additional Results for Real Data Analysis in Section LABEL:sec:realdata
We list the coefficients of the ITR $\widehat{f}_{\text{QIDS}}$ of the primary outcome estimated using SepL, FITR-IntL and FITR-Ramp in Table A.4, along with the secondary outcome ITRs $\widetilde{f}_{\text{CGI}}$ and $\widetilde{f}_{\text{SAS}}$ estimated using SepL. They are learned with the complete dataset and the linear kernel. We can conclude that the coefficients fitted by different methods are generally close, which suggests that the fusion penalty will not dramatically change an ITR compared to SepL. The coefficients are more similar for QIDS-change and CGI, which is expected since SAS measures the impacts of depression on social functioning. The Flanker variable has a different sign in the ITR of SAS than the other two variables.
@@ B.1 Additional Assumptions and Theoretical Results
To bound the risk with 0-1 loss using the risk with convex surrogate loss, we need another common assumption in classification literature, which assumes the noise conditions around the decision boundary (Bartlett et al.,, 2006). For Assumption B.4, it has been shown that the boundary assumption of $\eta(\bm{x})$ regarding $\beta$ is equivalent to the misclassification assumption of $f_{1}$ regarding $\alpha$ (Bartlett et al.,, 2006).
@@
B.2 Additional Details for Section LABEL:sec:theory
The parameters $\lambda_{1n},\mu_{1n},\kappa_{1n}$ should balance the estimation error and approximation error to achieve the minimal upper bound $\delta_{1n}(\tau)$. The minimal approximation error is $\left(\mu_{1n}\kappa_{1n}^{-1}\right)^{\frac{1}{1+q}}\lambda_{1n}^{\frac{q}{1+q}}$ when $\sigma_{1n}=\left(\mu_{1n}\kappa_{1n}^{-1}\lambda_{1n}^{-1}\right)^{\frac{1}{(1+q)d}}$. Then the best $\lambda_{1n}$ that balances the estimation error and the approximation error is $\lambda_{1n}=\left(\mu_{1n}\kappa_{1n}^{-1}\right)^{\frac{2(\Delta-1)+2(1+q)\nu}{3q+2\Delta+1}}n^{-\frac{1+q}{3q+2\Delta+1}}$. When $\mu_{1n}=0$, the convergence rate of the risk is
@@ B.2 Additional Details for Section LABEL:sec:theory
$\inf_{f_{1}\in\mathcal{H}}\left[\mathcal{R}(f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\right]-\mathcal{R}(f_{1}^{*})\lesssim\lambda_{1n}\sigma_{1n}^{d}+\left(1+\frac{\mu_{1n}}{\kappa_{1n}}\right)(2d)^{qd/2}\sigma_{1n}^{-qd}.$ (B.12)
@@
Since $r/p_{0}+\mu_{1n}\Omega_{12}\simeq M$, $\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}$ is bounded by $\sqrt{M/\lambda_{1n}}$, where $M:=r/p_{0}$. To this end, it suffices to consider the ball of radius $\sqrt{M/\lambda_{1n}}$. Therefore, the function class that we consider here is
@@
$\begin{split}&\log N\left(B^{-1}\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},\epsilon,L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}}),\frac{B\epsilon}{M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1}},L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}},\gamma_{n}\epsilon,L_{2}(\mathbb{P}_{n})\right),\end{split}$
where
$\gamma_{n}:=\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}}$
@@
$\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\leq\log N\left(B_{\mathcal{H}},\gamma_{n}\epsilon,L_{2}(\mathbb{P}_{n})\right)+\log\frac{M}{B\epsilon}.$
@@
$\sup_{\mathbb{P}_{n}}\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\lesssim\sigma_{1n}^{(1-\nu/2)(1+\delta)d}(\gamma_{n}\epsilon)^{-\nu},$
where $\sigma_{1n}>0$ is the parameter of the Gaussian kernel associated with $\mathcal{H}$, and $00<\nu\leq 2,\delta>0,\epsilon>$ 0. Therefore, we have $a=\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}$, $p=\nu$ and
@@
$\omega_{n}(\mathcal{G},\epsilon)\leq 2\operatorname{Rad}(\mathcal{G},n,\epsilon)\lesssim Ba^{\frac{1}{2}}n^{-\frac{1}{2}}\simeq\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}n^{-\frac{1}{2}}.$ (B.15)
@@
$\begin{split}&\left[\mathcal{R}(\widehat{f}_{1n})+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\right]-\left[\mathcal{R}(f_{1}^{\dagger})+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]\leq\sup_{h\in\mathcal{E}}\mathbb{P}_{n}h\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}n^{-\frac{1}{2}}+\sqrt{\frac{2\tau\lambda_{1n}^{-1}}{n}}+\frac{\tau\lambda_{1n}^{-\frac{1}{2}}}{n}\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}n^{-\frac{1}{2}}\left[\sqrt{\tau}+\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}\right]\end{split}$ (B.16)
@@
$\gamma_{n}:=\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}},$

### 09 Mar 2024 v3

@@
Appendix A ADDITIONAL DETAILS AND EXPERIMENTS FOR SECTIONS LABEL:sec:simulation AND LABEL:sec:realdata
In this section, we provide some details about implementation and experiments in the simulation study of Section LABEL:sec:simulation and the real data analysis of Section LABEL:sec:realdata.
@@ A.1 Implementation Details
A.2 Details for Data Generation and Value Function Estimation in Section LABEL:sec:simulation
@@
A.3 Additional Experiments for Estimating FITR in Section LABEL:sec:simulation
@@ A.5 Additional Results for Real Data Analysis in Section LABEL:sec:realdata
We list the coefficients of the ITR $\widehat{f}_{\text{QIDS}}$ of the primary outcome estimated using SepL, FITR-IntL and FITR-Ramp in Table A.4, along with the secondary outcome ITRs $\widetilde{f}_{\text{CGI}}$ and $\widetilde{f}_{\text{SAS}}$ estimated using SepL. They are learned with the complete dataset and the linear kernel. We can conclude that the coefficients fitted by different methods are generally close, which suggests that the fusion penalty will not dramatically change an ITR compared to SepL. The coefficients are more similar for QIDS-change and CGI, which is expected since SAS measures the impacts of depression on social functioning. The Flanker variable has a different sign in the ITR of SAS than in the other two outcomes.
@@ B.1 Additional Assumptions and Theoretical Results
To bound the risk with 0-1 loss using the risk with convex surrogate loss, we need another common assumption in classification literature, the Tsybakov’s noise assumption, which assumes the noise conditions around the decision boundary (Bartlett et al.,, 2006).
@@
For Assumption B.4, it has been shown that the boundary assumption of $\eta(\bm{x})$ regarding $\beta$ is equivalent to the misclassification assumption of $f_{1}$ regarding $\alpha$ (Bartlett et al.,, 2006).
@@
B.2 Analysis Details for Section LABEL:sec:theory
The parameters $\lambda_{1n},\mu_{1n},\kappa_{1n}$ should balance the estimation error and approximation error to achieve the minimal upper bound $\delta_{1n}(\tau)$. The minimal approximation error is $\gamma_{1n}^{\frac{1}{1+q}}\lambda_{1n}^{\frac{q}{1+q}}$ when $\sigma_{1n}=\left\{\gamma_{1n}\lambda_{1n}^{-1}\right\}^{\frac{1}{(1+q)d}}$. Then the best $\lambda_{1n}$ that balances the estimation error and the approximation error is $\lambda_{1n}=\gamma_{1n}^{\frac{2(\Delta-1)+2(1+q)\nu}{3q+2\Delta+1}}n^{-\frac{1+q}{3q+2\Delta+1}}$. When $\mu_{1n}=0$, the convergence rate of the risk is
@@ B.2 Analysis Details for Section LABEL:sec:theory
To compare the agreement rate with or without the fusion penalty, we break down equation (LABEL:equ:compare.SepL.FITR) as follows. When $\mu_{1n}\gtrsim\kappa_{1n}$, we have $\gamma_{1n}\simeq\mu_{1n}/\kappa_{1n}$. If $\alpha=1/2,\nu=3/2,\delta=1,\Delta=1/2$, the first condition in (LABEL:equ:compare.SepL.FITR) is satisfied with $\kappa_{1n}\gtrsim n^{-2q/(9q+6)},\mu_{1n}\gtrsim\kappa_{1n}$, and $N_{2}\gtrsim n^{3}\kappa_{1n}^{(9q+6)/q}\gtrsim n$. On the other hand, when $\mu_{1n}\lesssim\kappa_{1n}$, we have $\gamma_{1n}\simeq 1$. For the same $\alpha,\nu,\delta,\Delta$, the second condition in (LABEL:equ:compare.SepL.FITR) is satisfied with $\mu_{1n}\gtrsim n^{-2q/(9q+6)},\kappa_{1n}\gtrsim\mu_{1n}$, and $N_{2}\gtrsim n^{3}\mu_{1n}^{(9q+6)/q}\gtrsim n$.
@@
$\inf_{f_{1}\in\mathcal{H}}\left[\mathcal{R}(f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\right]-\mathcal{R}(f_{1}^{*})\lesssim\lambda_{1n}\sigma_{1n}^{d}+\gamma_{1n}(2d)^{qd/2}\sigma_{1n}^{-qd},$ (B.12)
where $\gamma_{1n}:=1+\mu_{1n}/\kappa_{1n}$.
@@
Since $r/p_{0}+\mu_{1n}\Omega_{12}\simeq M$ with $M:=r/p_{0}$, $\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}$ is bounded by $\sqrt{M/\lambda_{1n}}$. To this end, it suffices to consider the ball of radius $\sqrt{M/\lambda_{1n}}$. Therefore, the function class that we consider here is
@@
$\begin{split}&\log N\left(B^{-1}\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},\epsilon,L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}}),\frac{B\epsilon}{M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1}},L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}},\gamma_{1n}^{-1}\epsilon,L_{2}(\mathbb{P}_{n})\right),\end{split}$
since
$\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}}=\gamma_{1n}^{-1}.$
@@
$\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\leq\log N\left(B_{\mathcal{H}},\gamma_{1n}^{-1}\epsilon,L_{2}(\mathbb{P}_{n})\right)+\log\frac{M}{B\epsilon}.$
@@
$\sup_{\mathbb{P}_{n}}\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\lesssim\sigma_{1n}^{(1-\nu/2)(1+\delta)d}(\gamma_{1n}^{-1}\epsilon)^{-\nu},$
where $\sigma_{1n}>0$ is the parameter of the Gaussian kernel associated with $\mathcal{H}$, and $00<\nu\leq 2,\delta>0,\epsilon>$ 0. Therefore, we have $a=\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{1n}^{\nu}$, $p=\nu$ and
@@
$\omega_{n}(\mathcal{G},\epsilon)\leq 2\operatorname{Rad}(\mathcal{G},n,\epsilon)\lesssim Ba^{\frac{1}{2}}n^{-\frac{1}{2}}\simeq\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{1n}^{\nu}n^{-\frac{1}{2}}.$ (B.15)
@@
$\begin{split}&\left[\mathcal{R}(\widehat{f}_{1n})+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\right]-\left[\mathcal{R}(f_{1}^{\dagger})+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]\leq\sup_{h\in\mathcal{E}}\mathbb{P}_{n}h\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{1n}^{\nu}n^{-\frac{1}{2}}+\sqrt{\frac{2\tau\lambda_{1n}^{-1}}{n}}+\frac{\tau\lambda_{1n}^{-\frac{1}{2}}}{n}\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}n^{-\frac{1}{2}}\left[\sqrt{\tau}+\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{1n}^{\nu}\right]\end{split}$ (B.16)
@@
$\gamma_{1n}^{-1}:=\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}},$